# The three phases of self-gravitating scalar field ground states

### v2

@@ Abstract
It is generally assumed that scalar field dark matter halos would contain solitonic cores—spherically symmetric ground state configurations—at their centers. This is especially interesting in the case of ultralight dark matter (ULDM), where the solitons sizes are on the order of galaxies. In this work, we show that the paradigm of a spherically symmetric soliton embedded in the center of each halo is not universally valid in a scenario with multiple interacting scalar fields. In particular, sufficiently strong repulsive interspecies interactions make the fields immiscible. In such models, the ground state configuration can fall into a number of different phases that depend on the fields’ relative densities, masses, and interaction strengths. This raises the possibility that the inner regions of ULDM halos are more complex and diverse than previously assumed.
preprint: APS/123-QED
Introduction
Gravitating scalar fields are important components of theoretical models in cosmology. They arise naturally from certain overarching theories of quantum gravity and can have a variety of observational consequences in the Universe, including as ultralight dark matter (ULDM) Ferreira (2021), and in models of inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982); Martin et al. (2014). In this letter, we discuss the ground states and equilibria of one family of real scalar fields.
In scalar ULDM scenarios, the wave-like nature of the field becomes consequential on larger spatial scales set by the de Broglie wavelength, $\lambda_{\rm{dB}}\propto m^{-1}$. They form Bose-Einstein condensates on this scale. For a particle mass of $m\sim 10^{-19}-10^{-22}$, the de Broglie wavelength becomes evident on galactic scales ($\lambda_{\rm{dB}}\sim 1\,\rm{kpc}$) when the particle has no non-gravitational interactions Hu et al. (2000).
Figure 1: Illustrations of the three classes of ground state configurations discussed in this paper. Each panel shows a slice through the center of a 3D simulation box. Upper panels show the densities of the individual fields; lower panels show the corresponding total density. For comparison, single-field solitons of equivalent total and particle masses are shown in light gray. The existence of three distinct classes of ground states and the resulting diversity in halo density shapes is the main result of this work.
Most work on ULDM assumes that each dark matter halo has a solitonic core at its center.In ULDM, solitons condense quickly enough that most halos would contain a soliton Levkov et al. (2018); Eggemeier and Niemeyer (2019); Kirkpatrick et al. (2020); Chen et al. (2021); Hertzberg et al. (2020); Kirkpatrick et al. (2022); García et al. (2023). The role solitons could play as a solution to the cusp-core problem contributed to ULDM rising in popularity as a dark matter candidate Su and Chen (2011); Hui et al. (2017). While recent work has investigated the relative size and shape of solitonic cores in simulated ULDM halos Chan et al. (2022); Yavetz et al. (2022); Chan et al. (2022); Kendall et al. (2023); Zagorac et al. (2023); Dave and Goswami (2023, 2024); Chakrabarti et al. (2022), their existence in centers of halos has not been called into question. Nevertheless, the string axiverse generically predicts the existence of multiple ultralight fields with nonzero self- and interspecies interactions Arvanitaki et al. (2010); Arvanitaki and Dubovsky (2011), which in turn impacts the core profiles of such structures Glennon et al. (2023). Therefore, the mapping from interaction strengths to core profiles to the diversity of halos found in nature could enable constraints on axiverse models.
These results also have implications for inflation. In some inflationary models, the reheating epoch is analogous to structure formation in ULDM  Easther et al. (2011); Musoke et al. (2020), going as far as the formation of inflaton halos and solitonic inflaton stars Eggemeier and Niemeyer (2019). Interactions between multiple scalar inflaton fields would significantly change these collapsed states, impacting their dynamics during reheating and the consequent observational predictions Hotinli et al. (2018).
We present a detailed analysis of the equilibria of two-species models with interspecies interactions. We find that there is a wider variety of possible equilibria than previously known. In particular, we find that sufficiently repulsive interactions would imply immiscible phases, mimicking immiscible behavior that occurs in dual species Bose-Einstein condensates of ultracold atoms Ho and Shenoy (1996); Myatt et al. (1997); Esry et al. (1997); Timmermans (1998); Lee et al. (2016); Jiang et al. (2019); Gutierrez et al. (2021). Even under the assumption of spherical symmetry, sufficiently repulsive interactions imply the existence of ground states in which the maxima of the two species are not co-located. Upon relaxing the assumption of spherical symmetry, we find that ground states are not necessarily spherically symmetric. This picture would become increasingly complex with more than two species of scalar particles. This is, to our knowledge, the first instance of non-spherically symmetric ground states in cosmological scalar fields.
Background
Multispecies ULDM is governed by a system of coupled Gross-Pitaevskii-Poisson (GPP) equations:
@@ Introduction
$\nabla^{2}\Phi=4\pi G\sum_{j}m_{j}|\psi_{j}|^{2}$ (2)
@@
where $\Phi$ is the Newtonian gravitational potential, and the indices $j,k$ run over the $N$ fields considered, each with particle mass $m_{j}$. The self- and interspecies interactions are parameterized by a symmetric matrix $\lambda_{jk}$ with positive (negative) values corresponding to repulsive (attractive) interactions. The GPP equations arise from the Newtonian limit of the ULDM Lagrangian Glennon et al. (2023).
@@
$\rho_{j}(\mathbf{x},t)=m_{j}\left|\psi_{j}(\mathbf{x},t)\right|^{2},$ (3)
@@
The conserved energy of equations 1 and 2 is
$E$ $=E_{\text{grav}}+E_{\text{KQ}}+\sum_{j}E_{\text{self}}^{j}+\sum_{j}\sum_{k}E_{\text{int}}^{j,k}$ (4)
where the gravitational potential energy is
$E_{\text{grav}}=\int\mathrm{d}x^{3}\frac{1}{2}\Phi\sum_{j}m_{j}|\psi_{j}|^{2}\,.$ (5)
@@
The total kinetic energy consists of a classical term (arising from bulk motion) and quantum term. [^1] Note that the “quantum” kinetic energy does not have a quantum origin, but simply refers to the kinetic energy that does not correspond to any bulk motion Niemeyer (2019). Throughout this work we only consider equilibria, so the classical kinetic energy is zero and the total kinetic energy will be equal to the quantum kinetic energy:
$E_{\text{KQ}}=\frac{\hbar^{2}}{2}\int\mathrm{d}x^{3}\sum_{j}\frac{1}{m_{j}}\left|\nabla\psi_{j}\right|^{2}$ (6)
The energies due to self-interactions and inter-field interactions respectively are
$E_{\text{self}}^{j}=\frac{\hbar^{3}\lambda_{jj}}{2m_{j}^{2}}\int\mathrm{d}x^{3}|\psi_{j}|^{4};E_{\text{int}}^{j,k}=\frac{\hbar^{3}\lambda_{jk}}{2m_{j}^{2}}\int\mathrm{d}x^{3}|\psi_{k}|^{2}|\psi_{j}|^{2}\,.$ (7)
@@
In our investigations of the conditions of the system’s immiscibility—or separation of ground states, as presented in the rightmost panel of Fig. 1—we will make particular use of the energy definitions above. Our units and code used are defined in the Appendix.
Analytic evidence for phase transitions
@@
Before solving our system numerically, we present derivations of exact expressions for the energies under a spherically symmetric Gaussian ansatz for the ground state profiles. For single-species ULDM solitons, this ansatz has close agreement with numerical solutions Chavanis (2011); Glennon and Prescod-Weinstein (2021). We assume each field has the profile defined by its mass $M_{i}$ and length scale $R_{i}$, with centres of the profiles separated by a distance $d$:
$\rho_{i}(\mathbf{r})=\frac{M_{i}}{\pi^{3/2}R_{i}^{3}}\exp\left[-r^{2}/R_{i}^{2}\right].$ (8)
@@
With this ansatz, we can find analytic expressions for the quantum kinetic, gravitational, and interaction energies. As mentioned above, the quantum kinetic energy does not depend on the separation distance $d$, and therefore plays no role in determining which phase is preferred. The interaction energy can be calculated from Eq. (7),
$E_{\mathrm{int}}^{12}(d)$ $=\frac{\hbar^{3}\lambda_{12}}{4\mu^{2}}\frac{M_{1}M_{2}}{\pi^{3/2}(R_{1}^{2}+R_{2}^{2})^{3/2}}\exp\left[-\frac{d^{2}}{R_{1}^{2}+R_{2}^{2}}\right],$ (9)
@@
$E_{\mathrm{grav}}^{i}=\int d^{3}\mathbf{r}\rho_{i}(\mathbf{r})\Phi_{i}(\mathbf{r})=\frac{-1}{\sqrt{2\pi}}\frac{GM_{i}^{2}}{R_{i}},$ (10)
where $\Phi_{i}$ denotes the gravitational potential generated by species $i$. To calculate the interspecies gravitational energy $E_{\mathrm{grav}}^{12}$, we consider two cases. First when $d=0$, we can calculate it exactly in the Gaussian ansatz,
$E_{\mathrm{grav}}^{12}(d=0)=\int d^{3}\mathbf{r}\left(\rho_{1}\Phi_{2}+\rho_{2}\Phi_{1}\right)$ (11)
$=$ $\frac{2GM_{1}M_{2}}{\sqrt{\pi}}\left[\frac{1}{\sqrt{R_{1}^{2}+R_{2}^{2}}}-\frac{R_{1}^{2}}{(R_{1}^{2}+R_{2}^{2})^{\frac{3}{2}}}-\frac{R_{1}^{2}}{(R_{1}^{2}+R_{2}^{2})^{\frac{3}{2}}}\right].$ (12)
A simplification occurs when the two species have equal total mass $M$, particle mass $m$, and radii $R$. In this case the total gravitational energy evaluates to
$E_{\mathrm{grav}}(d=0)=\frac{-4}{\sqrt{2\pi}}\frac{GM^{2}}{R}.$ (13)
This agrees with the total gravitational energy of a single-species soliton in the Gaussian ansatz.
Second, at finite separation distance $d>0$, the interspecies gravitational energy $E_{\mathrm{grav}}^{12}$ must be evaluated numerically. However, it can be approximated at large separation distances $d\gg R_{1},R_{2}$ by noting that in this case the solitons become effectively point masses such that
$E_{\mathrm{grav}}^{12}(d\gg R_{i})=\frac{-GM_{1}M_{2}}{d}.$ (14)
The gravitational energy $E_{\mathrm{grav}}$ and the energy $E_{\mathrm{int}}$ due to repulsive interactions are in competition with each other: $E_{\mathrm{grav}}$ is minimized at $d=0$, while $E_{\mathrm{int}}$ is minimized at $d=\infty$. This suggests that there is a critical interaction strength at which the energy minimizing separation distance changes from $d=0$ to $d>0$. We can estimate the critical interaction strength $\lambda_{12}^{*}$ by comparing the total energy at $d=0$ to the total energy at $d=\infty$, $\Delta E=E(d=0)-E(d=\infty)$. Since the quantum kinetic energy and self-gravitational energies are independent of $d$ and both interspecies gravitational energy and interaction energy go to zero as $d$ becomes infinite, this simplifies to $\Delta E=E_{\mathrm{int}}(d=0)+E_{\mathrm{grav}}^{12}(d=0).$ When $\Delta E$ is positive, the system has lower total energy at infinite separation distance than at zero separation distance, indicating that the system is not in the nested miscible phase. Thus, solving $\Delta E=0$ gives an estimate for the critical interaction strength. When the two species have equal total mass and particle mass, this results in
$\lambda_{12}^{*}\approx\frac{8\pi GR^{2}m^{2}}{\hbar^{3}}.$ (15)
@@
$R$ $=\frac{3\sqrt{\pi}\hbar^{2}}{4\sqrt{2}GMm^{2}}\left(1+\sqrt{1+\frac{GM^{2}m^{2}\lambda_{12}}{6\pi^{2}\hbar}}\right).$ (16)
@@
Together with Eq. (15), this allows us to estimate the critical interaction strength for a given total mass. For two fields of particle mass $m=1$, and soliton mass $M=50$, this results in an estimated critical interaction strength of $\lambda_{12}^{*}=0.09$. This closely approximates the critical interaction strength we observe in numerical simulations presented below, where we find $\lambda_{12}^{*}=0.1$.
Numerical results and phase diagram
Having established an expectation for a critical interspecies interaction value $\lambda_{12}^{*}$ where a phase transition will occur, we now relax our Gaussian ansatz to find more realistic ground state profiles. For a single self-gravitating scalar field, the ground state profile is called a soliton. Solitons are static spherically symmetric states with balanced gravitational and “quantum”/gradient forces. The density profile of a single-field soliton without self-interactions is well-approximated as
$\rho(r)=\frac{\rho_{s}}{{\left(1+{(r/r_{s})}^{2}\right)}^{8}},$ (17)
where $\rho_{s}$ is the central density and the scale radius is
$r_{s}\simeq 0.335\,\frac{10^{9}M_{\odot}}{M}\left(\frac{10^{-22}\mathrm{eV}}{m}\right)^{2}\,\mathrm{kpc}\,.$ (18)
The total mass of this profile is
$M=2.2\times 10^{8}\left(\frac{m}{10^{-22}\mathrm{eV}}\right)^{-2}\left(\frac{r_{s}}{\mathrm{kpc}}\right)^{-1}M_{\odot}\,.$ (19)
The radius of a soliton decreases with both the mass $m$ of its constituent particles and its total mass $M$.
In realistic halos, the situation becomes more complicated. Solitons in power-law backgrounds are well approximated by changing the scale radius and exponent of $-8$ in equation 17 Guo (2020); Blum and Teodori (2021). Self-interactions in single field ULDM affect the possible types of equilibria and imply maximum soliton masses Chavanis (2023).

### v3

@@ Abstract
It is generally assumed that scalar field dark matter halos would contain solitonic cores—spherically symmetric ground state configurations—at their centers. This is especially interesting in the case of ultralight dark matter (ULDM), where the solitons sizes are on the order of galaxies. In this work, we show that the paradigm of a spherically symmetric soliton embedded in the center of each halo is not the only plausible structure formation scenario when there are multiple, interacting scalar fields. In particular, sufficiently strong repulsive interspecies interactions make the fields separate, creating the possibility of scenarios where there are distinct solitons. In such models, the ground state configuration can fall into a number of different phases that depend on the fields’ relative densities, masses, and interaction strengths. This raises the possibility that the inner regions of ULDM halos are more complex and diverse than previously assumed.
1 Introduction
Gravitating scalar fields are important components of theoretical models in cosmology: they arise naturally from certain overarching theories of quantum gravity and can have a variety of observational consequences in the Universe. In this paper, we discuss the ground states and equilibria of one family of real scalar fields. We are primarily interested in how these results can be applied to models of scalar dark matter (SDM) while also suggesting applications to models of inflation that include multiple scalar fields. The two applications are united in their reliance on the models’ ground states, which can inform how structures collapse and grow in both regimes, directly relating to the focus of this work. Our main result is that there three distinct classes of ground states in such theories given the existence of repulsive non-gravitational interactions, or couplings, between the fields. Examples of each class are shown in Fig. 1, and the roles gravitating scalar fields play in cosmology are outlined in Fig. 2. This is, to our knowledge, the first instance of non-spherically symmetric rotation-free ground states in cosmological scalar fields.
Both in studies of inflation and SDM, multifield models are not new. Indeed, models of inflation that include multiple scalar fields have been studied in some depth starting from the 1980s  [1, 2, 3, 4]. On the SDM side, a well-known motivation comes as a consequence of string theory: the string axiverse generically predicts the existence of multiple ultralight axion scalar fields with nonzero self- and interspecies interactions [5, 6]. We are therefore particularly interested in overlapping versions of ultralight dark matter (ULDM)[7]. There are several names for this class of models in the literature, and we take this opportunity to disambiguate our usage of the terminology and the assumptions behind each acronym in the following paragraphs. We also present the relationship between these terms graphically in the form of a Venn diagram embedded in Fig. 2.
@@ 1 Introduction
These Bose-Einstein Condensate Dark Matter (BECDM) models feature a Bose-Einstein condensate-like core often called a soliton. In single-field scenarios, the soliton corresponds to the ground state of the system and has a well known profile [9, 10, 11], though the exact shape is modulated by the presence of self-interactions [12]. The size of the soliton is inversely proportional to the mass of the constituent boson $m$; thus, different models admit solitons on different scales. Because a soliton is essentially the smallest collapsed structure that can be formed in a single-boson fielad, this means different models build or modify structure on different scales: for instance, stellar-sized solitons are sometimes referred to as Bose stars [13, 14]. For a particle mass of $m\sim 10^{-19}-10^{-22}$, the de Broglie wavelength becomes evident on galactic scales ($\lambda_{\mathrm{dB}}\sim 1\,\mathrm{kpc}$) [15]; this specific model is often referred to as fuzzy dark matter (FDM) and assumes the particle has no non-gravitational interactions. If such interactions are permitted, the model has been referred to as ultralight axions (ULAs). Finally, while we began our discussion with scalar fields, there is also overlap with some models of vector dark matter (VDM); we specifically refer to cases with multiple species of massive particles with integer spin $s$, mathematically equivalent to $2s+1$ scalar fields [16, 17].
Figure 1: Illustrations of the three classes of ground state configurations discussed in this paper. Each panel shows a slice through the center of a 3D numerical box. Each case has the same ratio of the two species’ axion masses ($m_{1}$ and $m_{2}$, where $m_{2}=10^{-22}\,\mathrm{eV}$), but the inter-species interaction strength ($\Lambda_{ij}$, presented in code units described in Appendix A) increases from left to right. Upper panels show the densities of the individual fields; lower panels show the corresponding total density. For comparison, single-field solitons of equivalent total and particle masses are shown in light gray. The existence of three distinct classes of ground states and the resulting diversity in halo density shapes is the main result of this work.
Figure 2: Scalar fields appear in many areas of cosmology, including as answers to the open questions of inflation, dark matter, and dark energy. We will particularly concern ourselves with their role as dark matter consisting of two ultralight axion-like particles (ULAs), though our results easily generalize to other applications. Considering how various scalar fields’ phenomenologies impact astrophysical observables could help narrow down the parameter space of theories of quantum gravity.
While our work presented here is generically applicable for many of the above mentioned models, we present results centered on parameter choices relevant to the “fuzzy” regime, where ground states or solitons correspond in size to galaxy cores. Most work on single-field ULDM, even when not analyzed as a Bose-Einstein condensate, assumes that each dark matter halo has a solitonic core at its center. In ULDM, solitons condense quickly enough that most halos would contain a soliton [14, 18, 19, 20, 21, 22, 23]. Indeed, the role solitons could play as a solution to the cusp-core problem contributed to ULDM rising in popularity as a dark matter candidate [24, 25]. It is generally understood that solitons exist in halo cores and are roughly spherical, although recent work has clarified that there is some diversity in how they relate to underlying haloes [26, 27, 26, 28, 29, 30, 31, 32]. Furthermore, while solitons are a feature of ULDM halos, it is important to emphasize that they are not a stationary structures and do exhibit oscillations, fluctuations, and motions within the halo core (see e.g. [33]); thus, while understanding the basic soliton structures of a model are a crucial first step, their behavior in vacuum is not fully representative of their behavior in a realistic dark matter halo.
Moving beyond single-field dark matter models, studies of ULDM have begun to consider how solitonic structures would look in multifield scenarios such as the axiverse. Changing the number of dark matter fields and allowing for non-gravitational interactions amongst them certainly impacts the ground states (and therefore the core profiles) of such structures [34, 35, 36, 37, 38, 39, 40]. The rising levels of interest in such axiverse scenarios are supported in the recent literature [37, 38, 41, 42, 43], including a particular emphasis on two-field scenarios and their core properties [44, 45, 40]. Thoroughly understanding the predictions of the simplest multi-field scenario is crucial for building intuition for more complicated axiverse predictions. Indeed, comparing qualitative differences in predicted ground state profiles with the latest data could help us constrain huge swathes of the axiverse parameter space, thus focusing our attention on the most astrophysically viable and interesting cases. Thus, it is valuable to understand the mapping between hypothetical interaction strengths and the ground states and the core profiles in which they ultimately result. A comparison between these results and the diversity of halos found in nature could constrain axiverse models, and even other other non-axionic dark sectors [46].
These same results also have implications for inflation. In some inflationary models, the reheating epoch is analogous to structure formation in ULDM [47, 48], going as far as the formation of inflaton halos and solitonic inflaton stars [18]. Interactions between multiple scalar inflaton fields would significantly change these collapsed states, impacting their dynamics during reheating and the consequent observational predictions [49]. Each of these applications will, of course, be highly model-dependent.
In this work, we present a detailed exploration of the equilibria of two-species models with quartic interspecies interactions, surveying a parameter space of interaction strengths and masses consistent with constraints from the string theory axiverse [50, 51]. We find that there is a wider variety of possible equilibria than previously known: in particular, sufficiently repulsive interactions can result in the separation of the two fields. This separation of the two fields mimics behavior that occurs in dual species Bose-Einstein condensates of ultracold atoms, which can have mixed (or “miscible”) and separated (or “immiscible”) phases [52, 53, 54, 55, 56, 57, 58]. Even when numerically imposing the assumption of spherical symmetry on the system, sufficiently repulsive interactions imply the existence of ground states in which the maxima of the two species are not co-located, i.e., with a density maximum in a shell a nonzero radius from the center (see middle column in Fig. 1). Upon relaxing the assumption, we find that ground states do not necessarily maintain spherical symmetry on their own and instead favor axisymmetric configurations (see right column in Fig. 1). This picture would become increasingly complex with more than two species of scalar particles; nevertheless, this asphericity presents new avenues for constraining this interesting but enormous parameter space.
2 Analytically Evaluating Multispecies ULDM for Phase Transitions
2.1 The System of Equations and Relevant Energies
We begin by concretizing the ULDM model we consider here. In general, multispecies ULDM is governed by a system of coupled Gross-Pitaevskii-Poisson (GPP) equations:
$i\hbar\frac{\partial\psi_{j}}{\partial t}=-\frac{\hbar^{2}}{2m_{j}}\nabla^{2}\psi_{j}+m_{j}\Phi\psi_{j}+\frac{\hbar^{3}}{2m_{j}^{2}}\lambda_{jj}|\psi_{j}|^{2}\psi_{j}+\frac{\hbar^{3}}{4m_{j}^{2}}\sum_{k\neq j}\lambda_{jk}|\psi_{k}|^{2}\psi_{j}$ (2.1)
$\nabla^{2}\Phi=4\pi G\sum_{j}m_{j}|\psi_{j}|^{2}$ (2.2)
@@ 2.1 The System of Equations and Relevant Energies
where $\Phi$ is the Newtonian gravitational potential, and the indices $j,k$ run over the $N$ fields considered, each with particle mass $m_{j}$. The self- and interspecies interactions are parameterized by a symmetric matrix $\lambda_{jk}$ with positive (negative) values corresponding to repulsive (attractive) interactions. The GPP equations arise from the Newtonian limit of the ULDM Lagrangian [38].
@@
$\rho_{j}(\mathbf{x},t)=m_{j}\left|\psi_{j}(\mathbf{x},t)\right|^{2},$ (2.3)
@@
The conserved energy of equations 2.1 and 2.2 can be written as
$E$ $=E_{\text{grav}}+E_{\text{KQ}}+\sum_{j}E_{\text{self}}^{j}+\sum_{j}\sum_{k>j}E_{\text{int}}^{j,k}\,.$ (2.4)
In order, there four energies refer to the gravitational potential energy, the kinetic energy, the energy of self-interactions, and the energy of interspecies interactions. We further define each type below, and they will be crucial to the arguments we make in subsequent sections.
The gravitational potential energy is
$E_{\text{grav}}=\int\mathrm{d}x^{3}\frac{1}{2}\Phi\sum_{j}m_{j}|\psi_{j}|^{2}\,.$ (2.5)
@@
The total kinetic energy consists of a classical term (arising from bulk motion) and quantum term. [^3] Note that the “quantum” kinetic energy does not have a quantum origin, but simply refers to the kinetic energy that does not correspond to any bulk motion [59]. Throughout this work we only consider equilibria, so the classical kinetic energy is zero and the total kinetic energy will be equal to the quantum kinetic energy:
$E_{\text{KQ}}=\frac{\hbar^{2}}{2}\int\mathrm{d}x^{3}\sum_{j}\frac{1}{m_{j}}\left|\nabla\psi_{j}\right|^{2}$ (2.6)
The energies due to self-interactions and interfield interactions respectively are
$E_{\text{self}}^{j}$ $=\frac{\hbar^{3}\lambda_{jj}}{2m_{j}^{2}}\int\mathrm{d}x^{3}|\psi_{j}|^{4}$ (2.7)
$E_{\text{int}}^{j,k}$ $=\frac{\hbar^{3}\lambda_{jk}}{2m_{j}^{2}}\int\mathrm{d}x^{3}|\psi_{k}|^{2}|\psi_{j}|^{2}\,.$ (2.8)
@@
In our investigations of the conditions of the system’s immiscibility—or separation of ground states, as presented in the rightmost panel of Fig. 1—we will make particular use of the energy definitions above.
2.2 Parameter choices and existing constraints
Before diving into the results of this work, we briefly comment on the parameters we center our exploration around in the context of existing constraints. Our results are present in scalable code units, discussed in more detail in Appendix A.
Throughout this work, we center our exploration on a particle mass of $m_{1}=m_{0}\equiv 10^{-22}\,\mathrm{eV/c^{2}}$ in one field, with the mass of the other parametrized as a ratio $m_{2}/m_{1}$. In the FDM scenario with no self-interactions, this region is heavily constrained [7] [^4] A useful summary of single-field limits can also be found on the website of Dr. Keir Rodgers at https://keirkwame.github.io/DM_limits/.. Given that it is the canonically “fuzzy” regime, we begin our investigation in the same place, with a view to eventually evaluating which mass constraints may hold in the interacting multifield regime.
The effect of self-interactions on ULAs has been explored elsewhere in the literature; they change the profile of ground state solitons, but do not fundamentally change their shape [38, 36, 35, 44]. Nevertheless, adding self-interactions to the single field model imposes certain limits: Ref. [60] derived limits on mass and self-interaction strengths below the fuzzy regime by comparing pressure from self-interaction to the quantum pressure, while Ref. [61] derives a constraint from the CMB power spectrum,
$\log\lambda<-91.86+4\log\left(\frac{m}{m_{0}}\right),$ (2.9)
which in our normalized units works out to $\Lambda\lesssim 3\Lambda_{0}$.
Inter-species interactions haven’t been subject to many astrophysically-motivated dark matter constraints thus far. Certain models of string theory can provide guidance: for instance, in Ref. [50] the authors use black hole superradiance to predict inter-species interactions are not expected to exceed the self-interaction strengths [50]. Given they are the most unconstrained feature of the model we explore, we were motivated to focus on exploration on how they affect potential observables. Consequently, we chose to put aside the effects of self-interactions by setting $\lambda_{ii}=0$ and only varying $\lambda_{ij}$. This allowed us to explore the effects of growing inter-species interactions without adding to an already high-dimensional space of variables; the detailed exploration of the effects of self-interactions we leave for future work.
2.3 Analytic evidence for phase transitions
In both the single and multi-field scenarios, the GPP equations Eq. (2.1) admit no closed form solutions. Numerical solutions for the ground state profile of a single self-gravitating scalar field are well studied [62, 63, 64, 65]. They are static spherically symmetric states with balanced gravitational and “quantum”/gradient forces, and are generally called solitons in the literature. Though exact closed form solutions do not exist, there are a number of well-studied approximations to the ground state soliton solution that can be useful for gaining physical insight and intuition. For a single self-gravitating field, the most robust ground state approximation is a power-law profile which has a high agreement to simulations [10],
$\rho(r)=\frac{\rho_{s}}{{\left(1+{(r/r_{s})}^{2}\right)}^{8}},$ (2.10)
where $\rho_{s}$ is the central density and the scale radius $r_{s}$ is
$r_{s}\simeq 0.335\,\frac{10^{9}M_{\odot}}{M}\left(\frac{10^{-22}\mathrm{eV}}{m}\right)^{2}\,\mathrm{kpc}\,.$ (2.11)
The total mass of this profile is
$M=2.2\times 10^{8}\left(\frac{m}{10^{-22}\mathrm{eV}}\right)^{-2}\left(\frac{r_{s}}{\mathrm{kpc}}\right)^{-1}M_{\odot}\,.$ (2.12)
Note that the radius of a soliton decreases with both the mass $m$ of its constituent particles and its total mass $M$. The above equations refer to the minimal FDM model; the addition of self-interactions affects the possible types of equilibria and imply maximum soliton masses [66].
The profile of non-isolated ground states is expected to be more complicated than the single-field soliton approximated by equation 2.10, which is conveniently universal when scaled by particle and soliton mass, $m$ and $M$. As with single-field modes, the GPP equations Eq. (2.1) governing multifield scenarios admit no closed form solutions. Furthermore, each field $\psi_{j}$ depends on the other fields $\psi_{k\neq j}$, already suggesting that universality of ground state profiles may be broken. We begin to build intuition about these states with energy minimization arguments using an analytic approximation for each field’s profile. The power law form of the above approximation makes using it for calculating derived quantities such as energies unwieldy, thereby losing the advantage of the analytic approach. For our purposes, a Gaussian profile is sufficient for approximating isolated solitons, and is useful for producing clean analytic estimates.
@@ 2.3 Analytic evidence for phase transitions
Thus, before solving our system numerically, we present derivations of exact expressions for the energies under a spherically symmetric Gaussian ansatz for the ground state profiles. For single-species ULDM solitons, this ansatz has close agreement with numerical solutions [12, 65]. We assume each field has the profile defined by its mass $M_{i}$ and length scale $R_{i}$, with centres of the profiles separated by a distance $d$:
$\rho_{i}(\mathbf{r})=\frac{M_{i}}{\pi^{3/2}R_{i}^{3}}\exp\left[-r^{2}/R_{i}^{2}\right].$ (2.13)
@@
With this ansatz, we can find analytic expressions for the quantum kinetic, gravitational, and interaction energies. As mentioned above, the quantum kinetic energy does not depend on the separation distance $d$, and therefore plays no role in determining which phase is preferred. The interaction energy can be calculated from Eq. (2.8),
$E_{\mathrm{int}}^{12}(d)$ $=\frac{\hbar^{3}\lambda_{12}}{4\mu^{4}}\frac{M_{1}M_{2}}{\pi^{3/2}(R_{1}^{2}+R_{2}^{2})^{3/2}}\exp\left[-\frac{d^{2}}{R_{1}^{2}+R_{2}^{2}}\right],$ (2.14)
@@
$E_{\mathrm{grav}}^{i}=\int d^{3}\mathbf{r}\rho_{i}(\mathbf{r})\Phi_{i}(\mathbf{r})=\frac{-1}{\sqrt{2\pi}}\frac{GM_{i}^{2}}{R_{i}},$ (2.15)
where $\Phi_{i}$ denotes the gravitational potential generated by species $i$.
To calculate the interspecies gravitational energy $E_{\mathrm{grav}}^{12}$, we consider two cases. First, when $d=0$ we can calculate it exactly in the Gaussian ansatz,
$E_{\mathrm{grav}}^{12}(d=0)$ $=\int d^{3}\mathbf{r}\left(\rho_{1}\Phi_{2}+\rho_{2}\Phi_{1}\right)$ (2.16)
$=\frac{2GM_{1}M_{2}}{\sqrt{\pi}}\left[\frac{1}{\sqrt{R_{1}^{2}+R_{2}^{2}}}-\frac{R_{1}^{2}}{(R_{1}^{2}+R_{2}^{2})^{\frac{3}{2}}}-\frac{R_{1}^{2}}{(R_{1}^{2}+R_{2}^{2})^{\frac{3}{2}}}\right].$ (2.17)
A simplification occurs when the two species have equal total mass $M$, particle mass $m$, and radii $R$. In this case, the total gravitational energy evaluates to
$E_{\mathrm{grav}}(d=0)=\frac{-4}{\sqrt{2\pi}}\frac{GM^{2}}{R}.$ (2.18)
This agrees with the total gravitational energy of a single-species soliton in the Gaussian ansatz. Second, at finite separation distance $d>0$, the interspecies gravitational energy $E_{\mathrm{grav}}^{12}$ must be evaluated numerically. However, it can be approximated at large separation distances $d\gg R_{1},R_{2}$ by noting that in this case the solitons become effectively point masses such that
$E_{\mathrm{grav}}^{12}(d\gg R_{i})=\frac{-GM_{1}M_{2}}{d}.$ (2.19)
The gravitational energy $E_{\mathrm{grav}}$ and the energy $E_{\mathrm{int}}$ due to repulsive interactions are in competition with each other: $E_{\mathrm{grav}}$ is minimized at $d=0$, while $E_{\mathrm{int}}$ is minimized at $d=\infty$. This suggests that there is a critical interaction strength at which the energy minimizing separation distance changes from $d=0$ to $d>0$. We can estimate the critical interaction strength $\lambda_{12}^{*}$ by comparing the total energy at $d=0$ to the total energy at $d=\infty$, $\Delta E=E(d=0)-E(d=\infty)$. Since the quantum kinetic energy and self-gravitational energies are independent of $d$, and both interspecies gravitational energy and interaction energy go to zero as $d$ becomes infinite, this simplifies to
$\Delta E=E_{\mathrm{int}}(d=0)+E_{\mathrm{grav}}^{12}(d=0).$ (2.20)
When $\Delta E$ is positive, the system has lower total energy at infinite separation distance than at zero separation distance, indicating that the system is not in the nested miscible phase, meaning that the two solitons will have started to separate. Thus, solving $\Delta E=0$ gives an estimate for the critical interaction strength. When the two species have equal total mass and particle mass, this results in
$\lambda_{12}^{*}\approx\frac{8\pi GR^{2}m^{4}}{\hbar^{3}}.$ (2.21)
@@
$R$ $=\frac{3\sqrt{\pi}\hbar^{2}}{4\sqrt{2}GMm^{2}}\left(1+\sqrt{1+\frac{GM^{2}m^{2}\lambda_{12}}{6\pi^{2}\hbar}}\right).$ (2.22)
@@
Together with Eq. (2.21), this allows us to estimate the critical interaction strength for a given total mass. For two fields of particle mass $m=m_{0}=10^{-22}$ eV, and soliton mass $M=50\mathcal{M}\approx 3.5\times 10^{9}{\left(10^{-23}\mathrm{eV}/m_{0}\right)}^{\frac{3}{2}}M_{\odot}$, this results in an estimated critical interaction strength of $\lambda_{12}^{*}=0.09\Lambda_{12}$, where the code unit $\Lambda_{12}$ is defined in the Appendix. This closely approximates the critical interaction strength we observe in numerical simulations presented below, where we find $\lambda_{12}^{*}=0.1\Lambda_{12}$.
3 Numerical results and phase diagram